# MambaIO: Global-Coordinate Inertial Odometry for Pedestrians via Multi-Scale Frequency-Decoupled Modeling

### v1

@@ Abstract
Inertial Odometry (IO) enables real-time localization using only acceleration and angular velocity measurements from an Inertial Measurement Unit (IMU), making it a promising solution for localization in consumer-grade applications. Traditionally, IMU measurements in IO have been processed under two coordinate system paradigms: the body coordinate frame and the global coordinate frame, with the latter being widely adopted. However, recent studies in drone scenarios have demonstrated that the body frame can significantly improve localization accuracy, prompting a re-evaluation of the suitability of the global frame for pedestrian IO. To address this issue, this paper systematically evaluates the effectiveness of the global coordinate frame in pedestrian IO through theoretical analysis, qualitative inspection, and quantitative experiments. Building upon these findings, we further propose MambaIO, which decomposes IMU measurements into high-frequency and low-frequency components using a Laplacian pyramid. The low-frequency component is processed by a Mamba architecture to extract implicit contextual motion cues, while the high-frequency component is handled by a convolutional structure to capture fine-grained local motion details. Experiments on multiple public datasets show that MambaIO substantially reduces localization error and achieves state-of-the-art (SOTA) performance. To the best of our knowledge, this is the first application of the Mamba architecture to the inertial odometry task.
@@ I Introduction
The Inertial Measurement Unit (IMU) serves as the primary data source for Inertial Odometry (IO), providing acceleration and angular velocity measurements[surveyILS]. Owing to its low cost, compact size, and minimal power consumption, the IMU is widely integrated into mobile devices such as smartphones, robots, and unmanned aerial vehicles (UAV).
IO aims to estimate the trajectory of a moving carrier (e.g., humans or drones) based solely on acceleration and angular velocity measurements from the IMU, enabling high-precision self-localization and tracking. Traditional IO approaches primarily rely on numerical integration, but noise and bias inherent in IMU measurements accumulate over time, leading to significant drift errors. In contrast, learning-based IO methods, trained end-to-end on large-scale IMU datasets, automatically uncover latent motion regularities and demonstrate superior performance in terms of localization accuracy, robustness, and cross-scenario generalizability[RoNIN].
Early learning-based IO methods attempted to estimate the pose of the carrier directly in the body coordinate frame. However, RoNIN reported that the body frame undergoes frequent changes during motion, resulting in discontinuous motion representations and associated issues. To mitigate these problems, RoNIN introduced a Heading-Agnostic Coordinate Frame (HACF), defined as any coordinate system whose $Z$-axis is aligned with the gravity direction, thereby providing a more stable and reliable motion representation. This approach has since become the prevailing data processing paradigm in the IO community. Nevertheless, recent studies in drone-based scenarios have shown that, under certain conditions, using the body frame directly can lead to substantially improved localization accuracy. This counterintuitive observation motivates a reassessment of a fundamental question:
Is the global coordinate representation truly beneficial for pedestrian IO?
To address this question, we first conduct a systematic analysis of the representational characteristics of inertial data under different coordinate frames. Through kinematic analysis and visualization, we verify that the global coordinate frame provides more discriminative feature representations for pedestrian motion. Building on this finding, we propose MambaIO, which decomposes IMU measurements into frequency components to enable targeted learning. Experimental results demonstrate that MambaIO further improves localization accuracy in pedestrian IO, with selected results shown in Fig. 3.
The main contributions of this work are summarized as follows:
  • Coordinate Frame Representation Analysis: We systematically compare the representational capabilities of the global and body-fixed coordinate frames and their impact on learning performance, empirically validating the necessity of adopting the global frame in pedestrian IO.
@@
  • Mamba Module: We introduce the Mamba architecture to specifically model the contextual motion information embedded in the low-frequency IMU measurements component.
  • Multi-Path Convolutional Module: We design a multi-scale convolutional module to extract local motion features from the high-frequency IMU measurements component, which collaborates with the Mamba module to achieve comprehensive modeling of motion dynamics.
@@ II Related Work
From a methodological perspective, research on IO can be broadly divided into two categories: traditional methods based on Newtonian mechanics and data-driven approaches leveraging machine learning.
@@
Newtonian Mechanics-Based Methods: Early approaches, exemplified by Strapdown Inertial Navigation Systems (SINS), estimate the position and orientation of a carrier by numerically integrating the acceleration and angular velocity outputs of IMUs. However, these methods are highly sensitive to measurement accuracy, as even minor noise and bias in IMU measurements accumulate over time during integration, leading to significant drift errors. To mitigate such errors, various physical priors have been introduced as corrective constraints. For instance, Pedestrian Dead Reckoning (PDR) assumes a stable gait cycle, while Zero Velocity Update (ZUPT) exploits the prior knowledge that the velocity of a foot-mounted IMU is zero during stance phases. Furthermore, multi-sensor fusion techniques have been widely adopted to enhance system robustness by incorporating auxiliary information such as gait phase detection, vision, LiDAR, or GPS signals. In these frameworks, IMU data are typically fused with external measurements through Extended Kalman Filters (EKF), Invariant Extended Kalman Filters (IEKF), or factor graph optimization for joint state estimation. Nevertheless, physics-based methods often rely on strong assumptions about motion patterns or require additional hardware support, which limits their generalization capability and deployment flexibility in complex real-world environments.
@@
Learning-Based Methods: With the rapid development of deep learning, researchers have explored end-to-end models that learn motion mappings from large-scale IMU datasets to improve the accuracy, robustness, and cross-scenario adaptability of IO. Early representative works such as RIDI[RIDI] and PDRNet[PDRNet] first classify the mounting position of IMU and then regress the human’s velocity or displacement. Subsequently, RoNIN identified the representation discontinuity inherent in the body-fixed frame used by RIDI—caused by changes in device orientation—and proposed the HACF, such as the global coordinate frame. RoNIN also systematically compared multiple network architectures, including TCN, LSTM, and ResNet. Following this work, most studies adopted global coordinate frame for IMU preprocessing, employing model architectures based on CNNs, LSTMs, Transformers, or their hybrid variants. For example, TLIO[TLIO] and LIDR[LIDR] extended RoNIN by integrating Stochastic Cloning Extended Kalman Filters (SCEKF) and Left Invariant Extended Kalman Filters (LIEKF), respectively; IMUNet[IMUNet], DeepILS[DeepILS], and DeepLite[Deep-Learning-for-Inertial-Positioning] focused on lightweight network design; CTIN[CTIN], SBIPTVT[SBIPTVT], and iMOT[iMOT] introduced attention mechanisms to enhance temporal modeling; RNIN-VIO, SCHNN, and SSHNN constructed multi-branch hybrid networks combining CNNs, LSTMs, and attention modules.
@@ III Methodology
To examine the influence of different coordinate frames on pedestrian IO and to explore potential improvements in localization accuracy, this section first presents formal definitions of the body frame and the global frame in Section III-A. Section III-B then provides qualitative evidence from three complementary analyses, demonstrating that—unlike findings reported in drone-based applications—the global frame yields more discriminative representations of pedestrian motion. Building on this observation, Section III-C introduces MambaIO, which operates in the global coordinate frame and decomposes IMU measurements using a Laplacian pyramid. Dedicated modules are subsequently designed to model the implicit information embedded in each frequency component, thereby substantially enhancing the localization accuracy of pedestrian IO.
@@ III-A Coordinate Frame Definitions
III-A1 Body Coordinate Frame
As shown in Fig.3, the IMU directly measures angular velocity and acceleration in its own body frame. These measurements serve as input data during training, while the corresponding labels (e.g., human velocity) are obtained from more accurate algorithms or external sensors (e.g., RoNIN mounted an additional smartphone on the chest to collect human velocity). For simplicity, sensor bias and noise are temporarily neglected. At time $t$, the measured quantities in the body frame are related to the true physical values as follows:
Input Data: $\left\{\begin{aligned} \boldsymbol{\omega}_{t,\text{meas}}^{\text{body}}&=\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}\\ \mathbf{a}_{t,\text{meas}}^{\text{body}}&=\mathbf{a}_{t,\text{true}}^{\text{body}}+\mathbf{R}_{t,\text{global}\to\text{body}}\,\mathbf{g}^{\text{global}}\end{aligned}\right$. (1)
@@ III-A1 Body Coordinate Frame
where $\boldsymbol{\omega}_{t,\text{meas}}^{\text{body}}$ and $\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}$ denote the measured and true angular velocities expressed in the body frame, respectively; $\mathbf{a}_{t,\text{meas}}^{\text{body}}$ and $\mathbf{a}_{t,\text{true}}^{\text{body}}$ denote the measured and true accelerations expressed in the body frame, respectively; $\mathbf{R}_{t,\text{global}\to\text{body}}$ is the rotation matrix that transforms vectors from the global frame to the body frame; and $\mathbf{g}^{\text{global}}=[0,0,g]^{\top}$ represents the gravitational acceleration vector in the global frame, with $g\approx 9.81\,\text{m/s}^{2}$. $\mathbf{v}_{t,\text{meas}}^{\text{body}}$ denotes the measured velocity of the carrier (e.g., human or UAV) expressed in the body frame, and $\mathbf{v}_{t,\text{true}}^{\text{global}}$ denotes the true velocity of the carrier expressed in the global frame.
III-A2 Global Coordinate Frame
The input data and labels in the global coordinate frame can be obtained by applying rotational transformations to the corresponding quantities in the body frame. Consequently, at time $t$, the model’s input data and labels are expressed as:
Input Data: $\left\{\begin{aligned} \boldsymbol{\omega}_{t,\text{meas}}^{\text{global}}&=\mathbf{R}_{t,\text{body}\to\text{global}}\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}\\ \mathbf{a}_{t,\text{meas}}^{\text{global}}&=\mathbf{R}_{t,\text{body}\to\text{global}}\mathbf{a}_{t,\text{true}}^{\text{body}}+\mathbf{g}^{\text{global}}\end{aligned}\right$. (2)
@@ III-A2 Global Coordinate Frame
Fundamentally, Equations (1) and (2) are mathematically equivalent, representing the same IMU and velocity measurements in different reference frames. However, as illustrated in Fig. 3, the IMU is typically not co-located with the center of mass of the carrier (e.g., a pedestrian or a drone), causing measurements expressed in the IMU body frame to inadequately capture the true motion dynamics of the carrier.
In drone-based scenarios, the IMU is rigidly attached to the carrier, and measurements can be accurately transformed to the carrier’s center of mass using a pre-calibrated fixed transformation matrix that encodes the relative translation and rotation between the IMU and the carrier. This process is straightforward and highly accurate; consequently, some studies argue that the linear coupling between IMU measurements and pose information in Equation (1) provides a sufficient learning signal.
In contrast, for Pedestrian IO, the transformation from the IMU to the human body’s center of mass varies significantly over time and is generally infeasible to obtain in real time. For example, in RoNIN, human velocity is estimated from a smartphone mounted on the chest, while IMU measurements may originate from another phone held arbitrarily by the user. Under such practical conditions, models often directly estimate the carrier’s true velocity from raw IMU readings, implicitly learning the complex, time-varying mapping between the IMU location and an approximate human center of mass.
In this context, the linear coupling described in Equation (1) is insufficient to fully characterize the non-rigid, time-varying nature of this mapping. By comparison, the nonlinear coupling embodied in Equation (2) offers a more appropriate representation of human motion. Moreover, the global coordinate frame explicitly aligns measurements to a consistent reference orientation, effectively mitigating disturbances caused by rapid changes in carrier heading and thereby yielding motion representations that are more consistent and semantically meaningful. This observation aligns with the findings reported in RoNIN, which attributes severe discontinuities in motion representation to the drastic frame-to-frame orientation variations inherent in the body-fixed frame.
In summary, under clearly defined coordinate systems and data collection protocols, using raw IMU measurements in the body frame to estimate carrier velocity is viable for drone IO. However, for pedestrian IO, the transformation between the IMU and the human center of mass is highly dynamic and difficult to model, leading to significant representational discontinuities—a phenomenon consistent with prior empirical observations. Fig. 4 further illustrates the fundamental differences between these two IO learning paradigms.
@@ III-B Representation Analysis
To systematically evaluate the impact of different coordinate frames on the feature representational capacity in pedestrian inertial odometry (IO) and to validate the analysis presented in Section III-A, we conduct comparative experiments using Principal Component Analysis (PCA), t-SNE, and trajectory inference visualization. All experiments are performed on six publicly available pedestrian inertial datasets. Key findings are summarized below; a complete analysis is provided in the appendix.
Figure 2: Comparison of metrics on the xxx dataset. Scatter points closer to the bottom-left corner indicate the algorithm has lower ATE and RTE, i.e., higher localization accuracy.
PCA constructs orthogonal principal components to replace correlated variables in high-dimensional data, thereby enabling dimensionality reduction and interpretable structure analysis. The analysis pipeline for inertial data under different coordinate frames is as follows: (1) concatenate the latent features extracted by the RoNIN-ResNet model from all test samples into a high-dimensional feature matrix; (2) apply PCA to this matrix and compute normalized singular values, each corresponding to the variance explained by one principal component; (3) calculate the cumulative explained variance (i.e., energy coverage ratio) of the top $k=50$ principal components to quantify the efficiency of capturing essential motion information. As shown in Fig. 3, under the RoNIN-ResNet model, the global coordinate frame (red curve) consistently achieves higher cumulative explained variance than the body-fixed frame (black curve), attaining greater energy coverage with the same number of principal components. This indicates that the global frame concentrates critical motion information more effectively—a trend consistently observed across all other datasets (see Appendix).
Figure 3: Comparison of metrics on the xxx dataset. Scatter points closer to the bottom-left corner indicate the algorithm has lower ATE and RTE, i.e., higher localization accuracy.
@@
t-SNE maps high-dimensional features into a two- or three-dimensional space while preserving local neighborhood relationships, making it suitable for qualitatively evaluating feature separability. We extract features from raw IMU data in both the body-fixed and global frames on the RoNIN dataset and apply t-SNE for dimensionality reduction, coloring each sample according to its ground-truth speed magnitude. As shown in Fig. 4, features in the body-fixed frame exhibit substantial mixing across different speed levels with indistinct boundaries, suggesting a limited ability to encode speed-related motion semantics. In contrast, features in the global frame naturally cluster by speed, forming well-separated, coherent regions that enhance feature separability and provide a more discriminative representation for downstream models. Consistent patterns are observed across all other datasets (see Appendix).
To further validate the influence of coordinate frame selection on practical localization performance, we train and evaluate the RoNIN-ResNet model using data represented in both coordinate frames, and compare the alignment between predicted and ground-truth trajectories. As shown in Fig. 5, trajectories inferred from the global frame exhibit superior smoothness and positional accuracy compared with those from the body-fixed frame, more faithfully reconstructing the true motion state. This observation is highly consistent with the preceding PCA and t-SNE analyses, demonstrating that enhanced feature representation directly translates into improved motion modeling capability. Additional trajectory visualizations across different network architectures (e.g., CNN, Transformer) and datasets are provided in the Appendix, all of which further corroborate this conclusion.
Synthesizing evidence from the three experimental perspectives, we derive the following key conclusions: the assumption of “linear superposition of motion information and pose variation” inherent in the body-fixed frame fails to capture the complex, non-rigid, and nonlinear relative motion between pedestrians and IMUs, resulting in structurally deficient feature representations. In contrast, the global coordinate frame substantially enhances both feature compactness and class separability, producing a more discriminative and consistent latent representation for pedestrian IO. This representational advantage ultimately leads to more accurate and robust trajectory estimation, thereby validating both the effectiveness and necessity of adopting the global coordinate frame in pedestrian inertial odometry.

### v2

@@ Abstract
Inertial Odometry (IO) enables real-time localization using only acceleration and angular velocity measurements from an Inertial Measurement Unit (IMU), making it a promising solution for localization in consumer-grade applications. Traditionally, researchers have routinely transformed IMU measurements into the global frame to obtain smoother motion representations. However, recent studies in drone scenarios have demonstrated that the body frame can significantly improve localization accuracy, prompting a re-evaluation of the suitability of the global frame for pedestrian IO. To address this issue, this paper systematically evaluates the effectiveness of the global frame in pedestrian IO through theoretical analysis, qualitative inspection, and quantitative experiments. Building upon these findings, we further propose MambaIO, which decomposes IMU measurements into high-frequency and low-frequency components using a Laplacian pyramid. The low-frequency component is processed by a Mamba architecture to extract implicit contextual motion cues, while the high-frequency component is handled by a convolutional structure to capture fine-grained local motion details. Experiments on multiple public datasets show that MambaIO substantially reduces localization error and achieves state-of-the-art (SOTA) performance. To the best of our knowledge, this is the first application of the Mamba architecture to the IO task.
@@ I Introduction
The Inertial Measurement Unit (IMU) serves as the primary data source for Inertial Odometry (IO), providing acceleration and angular velocity measurements[Tartan-IMU]. Owing to its low cost, compact size, and minimal power consumption, the IMU is widely integrated into mobile devices such as smartphones, robots, and unmanned aerial vehicles (UAV)[AirIO].
IO aims to estimate the trajectory of a moving carrier (e.g., pedestrian or UAV) based solely on acceleration and angular velocity measurements from the IMU, enabling high-precision self-localization and tracking[surveyILS]. Traditional IO approaches primarily rely on numerical integration, but noise and bias inherent in IMU measurements accumulate over time, leading to significant drift errors[SINS]. In contrast, learning-based IO methods, trained end-to-end on large-scale IMU datasets, automatically uncover latent motion regularities and demonstrate superior performance in terms of localization accuracy and robustness, and cross-scenario generalizability[RoNIN].
Figure 1: Performance comparison of algorithms on the RoNIN dataset in terms of ATE and RTE. Points closer to the lower-left corner indicate lower errors and thus higher localization accuracy.
Early learning-based IO methods attempted to estimate the trajectory of the carrier directly in the body frame[6-DOF-IO]. However, RoNIN reported that the body frame undergoes frequent changes during motion, resulting in discontinuous motion representations and associated issues. To mitigate these problems, RoNIN introduced a Heading-Agnostic Coordinate Frame (HACF), defined as any coordinate system whose $Z$-axis is aligned with the gravity direction, thereby providing a more stable and reliable motion representation[RoNIN]. This approach has since become the prevailing data processing paradigm in the IO community. Nevertheless, recent studies in drone-based scenarios have shown that, under certain conditions, using the body frame directly can lead to substantially improved localization accuracy[AirIO]. This counterintuitive observation motivates a reassessment of a fundamental question:
Is the global frame representation truly beneficial for pedestrian IO?
To address this question, we first conduct a systematic analysis of the representational characteristics of inertial data under different coordinate frames. Through kinematic analysis and visualization, we verify that the global frame provides more discriminative feature representations for pedestrian motion. Building on this finding, we propose MambaIO, which decomposes IMU measurements into frequency components to enable targeted learning. Experimental results demonstrate that MambaIO further improves localization accuracy in pedestrian IO, with selected results shown in Fig. 1. The main contributions of this work are summarized as follows:
  • Coordinate Frame Representation Analysis: We systematically analyze the differences in motion representation across coordinate frames, validating the necessity of adopting the global frame in pedestrian IO.
@@
  • Mamba Module: We introduce the Mamba architecture to specifically model the contextual motion information embedded in the low-frequency component.
  • Multi-Scale Convolutional Module: We design a multi-scale convolutional module to extract local motion features from the high-frequency component, which collaborates with the Mamba module to achieve comprehensive modeling of motion dynamics.
@@ II Related Work
From a methodological perspective, research on IO can be broadly divided into two categories: traditional methods based on Newtonian mechanics and data-driven approaches leveraging machine learning[SurveyofIndoorInertial].
@@
Newtonian Mechanics-Based Methods: Early approaches, exemplified by Strapdown Inertial Navigation Systems (SINS), estimate the position and orientation of a carrier by numerically integrating the acceleration and angular velocity outputs of IMUs[SINS]. However, these methods are highly sensitive to measurement accuracy, as even minor noise and bias in IMU measurements accumulate over time during integration, leading to significant drift errors[EqNIO]. To mitigate such errors, various physical priors have been introduced as corrective constraints. For instance, Pedestrian Dead Reckoning (PDR) assumes a stable gait cycle[PDR], while Zero Velocity Update (ZUPT) exploits the prior knowledge that the velocity of a foot-mounted IMU is zero during stance phases[AdaptiveThreshold-BasedZUPT]. Furthermore, multi-sensor fusion techniques have been widely adopted to enhance system robustness by incorporating auxiliary information such as gait phase detection[PDR], vision[A-visual-inertial-approach-to-human-gait-estimation], or LiDAR[FAST-LIVO2]. In these frameworks, IMU data are typically fused with external measurements through Extended Kalman Filters (EKF)[RNIN-VIO, MSCKF] for joint state estimation. Nevertheless, physics-based methods often rely on strong assumptions about motion patterns or require additional hardware support, which limits their generalization capability and deployment flexibility in complex real-world environments[RoNIN].
@@
Learning-Based Methods: With the rapid development of deep learning, researchers have explored end-to-end models that learn motion mappings from large-scale IMU datasets to improve the accuracy, robustness, and cross-scenario adaptability of IO[surveyILS]. Early representative works such as RIDI[RIDI] and PDRNet[PDRNet] first classify the mounting position of IMU and then regress the pedestrian’s velocity or displacement. Subsequently, RoNIN identified the representation discontinuity inherent in the body frame used by RIDI—caused by changes in device orientation—and proposed the HACF, such as the global frame[RoNIN]. RoNIN also systematically compared multiple network architectures, including TCN, LSTM, and ResNet. Following this work, most studies adopted global frame for IMU preprocessing, employing model architectures based on CNNs, LSTMs, Transformers, or their hybrid variants. For example, TLIO[TLIO] and LIDR[LIDR] extended RoNIN by integrating Stochastic Cloning Extended Kalman Filters (SCEKF) and Left Invariant Extended Kalman Filters (LIEKF), respectively; IMUNet[IMUNet], DeepILS[DeepILS], and DeepLite[Deep-Learning-for-Inertial-Positioning] focused on lightweight network design; CTIN[CTIN], SBIPTVT[SBIPTVT], and iMOT[iMOT] introduced attention mechanisms to enhance temporal modeling; RNIN-VIO[RNIN-VIO], SCHNN[SCHNN], and SSHNN[SSHNN] constructed multi-branch hybrid networks combining CNNs, LSTMs, and attention modules.
@@ III Methodology
This section first provides formal definitions of IMU measurements and their corresponding labels in learning-based IO. Subsequently, through kinematic analysis and visualization-based comparison, we validate the necessity of the global frame in pedestrian IO—a conclusion that contrasts with prior findings[AirIO] in drone-based scenarios. Building on this insight, we propose MambaIO, a novel framework designed to more accurately learn the implicit motion states information embedded in IMU measurements within the global frame, thereby further enhancing pedestrian IO performance.
@@ III-A Coordinate Frame Definitions
III-A1 body frame
As shown in Fig. 2, the IMU directly measures angular velocity and acceleration in its own body frame. These measurements serve as training data, while the corresponding labels (e.g., pedestrian or UAV velocity) are obtained from more accurate algorithms or external sensors (e.g., RoNIN mounted an additional smartphone on the chest to collect pedestrian velocity[RoNIN]). For simplicity, sensor bias and noise are temporarily neglected. At time $t$, the measured quantities in the body frame are related to the true physical values as follows:
Training Data: $\left\{\begin{aligned} \boldsymbol{\omega}_{t,\text{meas}}^{\text{body}}&=\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}\\ \mathbf{a}_{t,\text{meas}}^{\text{body}}&=\mathbf{a}_{t,\text{true}}^{\text{body}}+\mathbf{R}_{t,\text{global}\to\text{body}}\,\mathbf{g}^{\text{global}}\end{aligned}\right$. (1)
@@
where $\boldsymbol{\omega}_{t,\text{meas}}^{\text{body}}$ and $\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}$ denote the measured and true angular velocities expressed in the body frame, respectively; $\mathbf{a}_{t,\text{meas}}^{\text{body}}$ and $\mathbf{a}_{t,\text{true}}^{\text{body}}$ denote the measured and true accelerations expressed in the body frame, respectively; $\mathbf{R}_{t,\text{global}\to\text{body}}$ is the rotation matrix that transforms vectors from the global frame to the body frame; and $\mathbf{g}^{\text{global}}=[0,0,9.81]^{\top}$ represents the gravitational acceleration vector in the global frame. $\mathbf{v}_{t,\text{meas}}^{\text{body}}$ denotes the measured velocity of the carrier (e.g., pedestrian or UAV) expressed in the body frame, and $\mathbf{v}_{t,\text{true}}^{\text{global}}$ denotes the true velocity of the carrier expressed in the global frame.
Figure 3: Comparison of IO Learning Processes in body frame and global frame.
III-A2 global frame
The training data and labels in the global frame can be obtained by applying rotational transformations to the corresponding quantities in the body frame. Consequently, at time $t$, the model’s training data and labels are expressed as:
Training Data: $\left\{\begin{aligned} \boldsymbol{\omega}_{t,\text{meas}}^{\text{global}}&=\mathbf{R}_{t,\text{body}\to\text{global}}\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}\\ \mathbf{a}_{t,\text{meas}}^{\text{global}}&=\mathbf{R}_{t,\text{body}\to\text{global}}\mathbf{a}_{t,\text{true}}^{\text{body}}+\mathbf{g}^{\text{global}}\end{aligned}\right$. (2)
@@
Fundamentally, Equations (1) and (2) are mathematically equivalent, representing the same IMU and velocity measurements in different reference frames. However, as illustrated in Fig. 2, the IMU is typically not co-located with the center of mass of the carrier (e.g., pedestrian or UAV), causing measurements expressed in the IMU body frame to inadequately capture the true motion dynamics of the carrier.
In drone-based scenarios, the IMU is rigidly attached to the carrier, and measurements can be accurately transformed to the carrier’s center of mass using a pre-calibrated fixed transformation matrix that encodes the relative translation and rotation between the IMU and the carrier. This process is straightforward and highly accurate; consequently, some studies argue that the linear coupling between IMU measurements and pose information in Equation (1) provides a sufficient learning signal[AirIO].
In contrast, for Pedestrian IO, the transformation from the IMU to the pedestrian body’s center of mass varies significantly over time and is generally infeasible to obtain in real time. For example, in RoNIN, pedestrian velocity is estimated from a smartphone mounted on the chest, while IMU measurements may originate from another phone held arbitrarily by the user[RoNIN]. Under such practical conditions, models often directly estimate the carrier’s true velocity from raw IMU readings, implicitly learning the complex, time-varying mapping between the IMU location and an approximate pedestrian center of mass.
In this context, the linear coupling described in Equation (1) is insufficient to fully characterize the non-rigid, time-varying nature of this mapping. By comparison, the nonlinear coupling embodied in Equation (2) offers a more appropriate representation of pedestrian motion. Moreover, the global frame explicitly aligns measurements to a consistent reference orientation, effectively mitigating disturbances caused by rapid changes in carrier heading and thereby yielding motion representations that are more consistent and semantically meaningful[RoNIN]. This observation aligns with the findings reported in RoNIN[RoNIN], which attributes severe discontinuities in motion representation to the drastic frame-to-frame orientation variations inherent in the body frame.
In summary, under clearly defined coordinate systems and data collection protocols, using raw IMU measurements in the body frame to estimate carrier velocity is viable for drone IO. However, for pedestrian IO, the transformation between the IMU and the pedestrian center of mass is highly dynamic and difficult to model, leading to significant representational discontinuities—a phenomenon consistent with prior empirical observations. Fig. 3 further illustrates the fundamental differences between these two IO learning paradigms.
@@ III-B Representation Analysis
To systematically validate the analysis presented in Section III-A, this section employs t-SNE visualization and trajectory inference visualization to conduct a comparative analysis of motion representations under different coordinate frames.
@@
t-SNE maps high-dimensional features into a two- or three-dimensional space while preserving local neighborhood relationships, making it suitable for qualitative assessment of feature separability. In our experiments, we apply t-SNE to IMU measurements represented in both the body and global frames, coloring each sample according to its ground-truth speed magnitude. As shown in Fig. 4, features in the body frame exhibit significant overlap across different speed ranges, with blurred boundaries, indicating limited capacity to encode speed-related motion semantics. In contrast, features in the global frame naturally cluster by speed, forming distinct, well-separated, and structurally coherent regions that substantially enhance feature discriminability and provide more informative representations for downstream models. This consistent pattern is observed across all other datasets (see Appendix for details).
Trajectory Visualization: To intuitively assess the impact of coordinate frame representations on learning in IO, we plot the trajectories inferred by the same network using identical IMU data but processed in either the body or global frame. As shown in Fig. 5, the red curve corresponding to the global frame closely aligns with the green ground-truth trajectory, whereas the blue curve representing the body frame deviates more significantly. This indicates that trajectories inferred from the global frame exhibit higher smoothness and positional accuracy, enabling more faithful reconstruction of the actual motion. This observation is highly consistent with the preceding t-SNE analyses, demonstrating that enhanced feature representation directly translates into improved motion modeling performance. This consistent pattern is observed across all other datasets (see Appendix for details).
In summary, through kinematic analysis and multi-perspective visualization comparisons, we conclude that the global frame facilitates more accurate and robust trajectory estimation, thereby validating its effectiveness and necessity in pedestrian IO.
Figure 4: t-SNE visualization of features from the RIDI dataset, demonstrating that the global frame representation yields better separability of motion patterns.